OpenStudy (anonymous):

Solve: 2x^3+x^2+x=0 any help?

OpenStudy (zzr0ck3r):

factor out an x and what do you have?

OpenStudy (anonymous):

^

OpenStudy (zzr0ck3r):

you can help bloop

OpenStudy (anonymous):

I'll let you have this one.

OpenStudy (anonymous):

x(x^2+x+1)=0?

OpenStudy (anonymous):

Divide both sides of the equation by x, and you have a quadratic equation. Solve.

OpenStudy (zzr0ck3r):

keep the x, its a 0

OpenStudy (zzr0ck3r):

you will have 3 answers

OpenStudy (zzr0ck3r):

0 is one of them, now solve the quadratic

OpenStudy (anonymous):

zzr0ck3r made mention to 3 solutions. Keep in mind that this is a mathematical law: The degree of a polynomial is equal to its roots.

OpenStudy (anonymous):

So should my final answer be (x=0) (x=-1) (x=-1)? This answer just doesn't seem right though...

OpenStudy (zzr0ck3r):

\[0=2x^3+x^2+x=x(2x^2+x+1)\] ok so x =0 is one solution so lets look at \[2x^2+x+1=0\\2x^2+x=-1\\ 2(x^2+\frac{x}{2})=-1\\2(x+\frac{1}{4})^2=-1+2(\frac{1}{4})^2=-1+\frac{1}{8}=\frac{-7}{8}\\so\\2(x+\frac{1}{4})^2=\frac{-7}{8}\\x=\pm\sqrt{\frac{-7}{16}}-\frac{1}{4}=\pm\frac{\pm\sqrt{7}i-1}{4}\]

OpenStudy (anonymous):

COMPLEX ROOTS FTW.

OpenStudy (zzr0ck3r):

the method I used to solve the quadratic was "completing the square" you could have used \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

thanks guys much love!